Height =12m diameter=4m of a well. After 5 mins the pump stops working. To remove the remaining water a pump of 9.84hp is used. Will the second pump be able to remove the water?
1
Expert's answer
2020-02-10T09:15:52-0500

Pump power is given by formula

"P=\frac{A}{t} (1)"

Where A is the work, t is the time

The work is given by formula

"A=mgh=ƿVgh=ƿShgh=ƿ\frac{πd×d}{4}hgh (2)"

We put (2) in (1)

"P=\frac{ ƿ πd×d hgh }{4t} (3)"

In our case, we have ƿ=103, d=4, h=12, g=10, t=5×60

Using (3) we get

P=80.8 hp (1hp=746 watts)

The second pump will not be able to pump out water because its power is less than 80.8 hp

Answer

The second pump will not be able to pump out water